OpenStudy (hang254):

A technician experimentally determined the boiling point of octane to be 124.1 C. The actual boiling point of octane is 125.7 C. Calculate the error and the percent error.

OpenStudy (anonymous):

err = 125.7- 124.1. percent_error = err *100 / 125.7.